a GAS EXPANDS BY 1.2 LITRES AT A CONSTANT PRESSURE OF 2.5 BAR. DURING THE EXPANSION, 500j OF HEAT IS ADDED. FIND THE CHANGE IN THE INTERNAL ENERGY OF THE GAS
1
Expert's answer
2019-06-06T10:10:31-0400

according to the first law of thermodynamics

"\Delta U = Q - p \Delta V = 500 - 1.2\cdot2.5\cdot10^{2} = 200 
 (J)"